OpenStudy (anonymous):

Which point is not on the graph of 2x + 3y – z – 12 = 0? (1, 1, 7) (0, 4, 0) (6, 0, 0) (3, 3, 3)

OpenStudy (anonymous):

To solve this, plug each point into the equation and see which one gives you a statement that is not true - that point is not on the graph (1,1,7): 2(1) + 3(1) - 7 - 12 = 2 + 3 - 7 - 12 = -14 Since -14 ≠ 0, that point is not on the graph.